A transformer with 40 turns on the high voltage winding is used to step down the voltage  from 240V to 120V. Find the number of turns in the low voltage winding.


1
Expert's answer
2021-05-05T07:45:19-0400

Using the expression for the transformation ratio

"k=\frac{U_1}{U_2}=\frac{N_1}{N_2}" Where "U_1=240V,U_2=120V,N_1=40"

write down

"N_2=\frac{N_1 \cdot U_2}{U_1}=\frac{40 \cdot 120}{240}=\frac{40}{2}=20(turns )"